The length of a rectangle is 7yards more than twice its width. Find the dimensions of the rectangle if the area of the rectangle is 39 square yards.
Answer by ewatrrr(24785)

 
Hi,
Let x and (2x+7)represent the width and length respectively
Question states***
x(2x+7) = 39 yds^2
2x^2 + 7x - 39 = 0
(2x + 13)(x-3) = 0
x = 3 |throwing out the Negative solution
Width = 3 yds and length = 13 yds
CHECKING our Answer***
3yds*13yds = 39 yds^2